# Decoding Intelligence: A Framework for Certifying Knowledge Comprehension in LLMs

### v2

@@ 3.1 Specifying knowledge comprehension
(Multi-hop reasoning problem from ${\mathcal{G}}$). Consider any path $\Pi$ (Definition 3.1) of length $l$ in ${\mathcal{G}}$. A multi-hop reasoning problem ${\mathcal{Q}}$ for $\Pi$ is identifying the tail node $\Pi_{T}$, given an alias of its head node $\Pi_{H}$ and aliases of all edges from $\Pi_{H}$ to $\Pi_{T}$ in $\Pi$. Let $v_{\mathcal{A}}$ denote the corresponding aliases of node $v$ in ${\mathcal{G}}$. Let $\mathcal{D}$ be a function that samples a random alias from the given set of aliases.
${\mathcal{Q}}\coloneqq\mathcal{D}(\Pi_{H,\mathcal{A}})\xrightarrow[]{\mathcal{D}((\Pi_{H},\Pi[2])_{\mathcal{A}})}\dots\xrightarrow[]{\mathcal{D}((\Pi[-2],\Pi_{T})_{\mathcal{A}})}?$
${\mathcal{Q}}$ thus involves $l-1$ reasoning steps to get to the final answer, where each step requires correctly identifying intermediate nodes of $\Pi$. Note, however, that correctness of intermediate reasoning steps is generally not evaluated, and accuracy is defined for the final response (Rajpurkar et al., 2016).
@@
To aid ${\mathcal{L}}$ in correctly answering a multi-hop reasoning query ${\mathcal{Q}}$ and reduce hallucination (Dhuliawala et al., 2023), we provide relevant textual information needed to identify the intermediate and final nodes in the prompt. Hence, the overall task of answering ${\mathcal{Q}}$ involves information extraction for identifying intermediate nodes and reasoning to connect the intermediate answers to reach the final answer, which we collectively call knowledge comprehension. Our overall property quantifies the probability of observing correct knowledge comprehension for a random multi-hop reasoning query developed from ${\mathcal{G}}$. We formally define the property for ${\mathcal{L}}$ as a probabilistic program over ${\mathcal{G}}$ in Algorithm 1. We follow the syntax of the imperative probabilistic programming language in (Sankaranarayanan et al., 2013, Figure 3). The language has primitives for sampling from common distributions such as Uniform ($\mathcal{U}$), Bernoulli ($Ber$), etc., and an estimateProbability(.) function that outputs the probability of a random variable attaining a certain value. As all the random sampling steps in Algorithm 1 can operate with any discrete distribution, we use a generic identifier $\mathcal{D}$ (line 1) for samplers of discrete distributions ($\dots$’ denotes samplers for other discrete distributions).
[…]
A prompt for ${\mathcal{L}}$ consists of ${\mathcal{Q}}$ and a context $\Gamma$ containing information relevant to answer ${\mathcal{Q}}$. $\Gamma$ is formed by concatenating ($\odot$) the contexts for all nodes in $\Pi$. Let $v_{\gamma}$ denote the corresponding context of node $v$ in ${\mathcal{G}}$. Prior works (Shi et al., 2023) on analyzing reasoning in LMs have shown the negative influence of irrelevant information (distractor) in prompts on the performance of LMs, which is not ideally expected. Hence, we include distractor texts in $\Gamma$ and specify that the correct response should not be based on the distractor information. The contexts of nodes $\tilde{v}$ adjacent to any node $\Pi[i]$ ($i\in[1,l-2]$) on $\Pi$, such that the relation of $(\Pi[i],\tilde{v})$ is the same as that of $(\Pi[i],\Pi[i+1])$, can serve as effective distractors for ${\mathcal{Q}}$ (Definition 3.3). This is because, at any intermediate step, the model can pick $\tilde{v}$ as the response, which can deviate ${\mathcal{L}}$’s reasoning from $\Pi$. Nodes adjacent to $\Pi[-1]$ and $\Pi[-2]$ are not distractors. For the former, the model must have already reached the final answer before getting to its adjacent nodes, hence, answering ${\mathcal{Q}}$. In the latter, adjacent nodes following same relation are valid correct answers and not distractors. We denote distractor text in $\Gamma$ as the context of randomly sampled nodes from a distribution $\mathcal{D}$ over all distractor nodes of $\Pi$ in ${\mathcal{N}}$. We demonstrate the effects of using distractor text on the performance of SOTA LLMs in Section 4.
Input: ${\mathcal{L}},{\mathcal{G}},v_{1},\rho$
@@
2:  $\Pi\coloneqq[v_{1},v_{2}\coloneqq(\mathcal{D}([v^{\prime}\mid(v_{1},v^{\prime})\in{\mathcal{E}}])),\dots,v_{\mathcal{D}(\{2,\dots,\rho\})}\coloneqq(\mathcal{D}([v^{\prime}\mid(v_{k},v^{\prime})\in{\mathcal{E}}]))]$
3:  ${\mathcal{Q}}\coloneqq\mathcal{D}(\Pi[0]_{\mathcal{A}})\xrightarrow[]{\mathcal{D}(\Pi[0],\Pi[1])_{\mathcal{A}}}\dots\xrightarrow[]{\mathcal{D}(\Pi[-2],\Pi[-1])_{\mathcal{A}}}\textbf{?}$
4:  $\Gamma\coloneqq\texttt{shuffle}([\Pi[0]_{\gamma},\dots,\Pi[-1]_{\gamma},(\mathcal{D}({\mathcal{N}}))_{\gamma},\dots])$
5:  $\mathcal{P}\coloneqq\Gamma\odot{\mathcal{Q}}$
6:  $p\coloneqq\texttt{estimateProbability}({\mathcal{L}}(\mathcal{P})==\texttt{any}(\Pi[-1]_{\mathcal{A}}))$
Algorithm 1 Knowledge comprehension specification
@@
(Distractor node). A distractor node $\tilde{v}$ for a path $\Pi=[v_{1},v_{2},\dots,v_{l}]$ of ${\mathcal{G}}$ is such that $\forall i\in[1,l],\tilde{v}\neq v_{i}$, and $\exists j\in[1,l-2],[(v_{j},\tilde{v})\in{\mathcal{E}}]\wedge[(v_{j},\tilde{v})_{\mathcal{A}}=(v_{j},v_{j+1})_{\mathcal{A}}]$.
Prior works such as (Chen et al., 2024) have shown that LLM performance can vary with information ordering. Hence, we shuffle the information in $\Gamma$ (line 4) to specify that the model’s response should be invariant to the ordering of information. Our final specification in line 6 is the probability that ${\mathcal{L}}$ generates any alias of the last node of the path, which is the correct answer to ${\mathcal{Q}}$. The specification depends on the choices for the different distributions used at various sampling steps, ${\mathcal{G}}$, $v_{1}$, and $\rho$. It leads to certificates for the behavior of ${\mathcal{L}}$ on a given subgraph of ${\mathcal{G}}$ and paths of length at most $\rho$.
@@ 3.2 Certification method
Our algorithm certifies the target LLM ${\mathcal{L}}$ by computing an interval $[p_{l},p_{u}]$ containing the value of the probability $p$ (Algorithm 1, line 6) for a given pivot node $v_{1}$ in ${\mathcal{G}}$ with high confidence. We model $p$ as the probability of setting the underlying boolean random variable $\mathcal{R}=({\mathcal{L}}(\mathcal{P})==\texttt{any}(\Pi[-1]_{\mathcal{A}}))$ to true (success). Thus, $\mathcal{R}\sim Ber(p)$. Exactly determining $p$ would require enumerating over all possible $\mathcal{P}$ which can be developed from any path from a subgraph of ${\mathcal{G}}$ with any random aliases, resulting in an infeasible number of possible prompts, as shown in Appendix B.6. Moreover, we want our method to generalize to closed-source LLMs as well, where the internal structures of the models are unknown. Hence, we cannot use any symbolic methods (Mirman et al., 2020) to determine $p$. Thus, to scalably certify the black-box target LM ${\mathcal{L}}$, we estimate $p$ with provably high-confidence (low error) bounds. Confidence is defined as the probability of the true value $p$ being within the bounds, i.e., $Pr[p\in[p_{l},p_{u}]]$. To establish formal guarantees, we want our estimation procedure to be such that the actual confidence is at least the user-specified confidence level, $1-\delta$ (i.e., $Pr[p_{l}\leq p\leq p_{u}]\geq 1-\delta$), where $\delta$ is a small positive constant. Hence we use conservative method of the Clopper-Pearson confidence intervals (Clopper & Pearson, 1934; Brown et al., 2001; Kurz et al., 2014; Collins, 2010), which is known to produce intervals that are guaranteed to have high confidence. To compute the high-confidence bounds on $p$, we make $n$ independent and identically distributed observations of $\mathcal{R}$, out of which we obtain $0\leq k\leq n$ successes. With the $n$ observations and $1-\delta$ confidence, we use Clopper-Pearson confidence intervals to bound $p$.
@@ 4 Experiments
4.1 Experimental setup
We certify the following open-source, instruction finetuned (Wei et al., 2022a) models — Llama-3-instruct 8B model (Dubey et al., 2024), Mistral 7B-Instruct-v0.2 (Jiang et al., 2023a), Phi-3 3B and 14B parameter models (Abdin et al., 2024). We also certify 4-bit and 8-bit quantized versions of the open-source models to study the effects of quantization on a model’s knowledge comprehension capabilities. Among the closed-source models with API access, we certify Gemini-1.5 (Gemini Team, 2024) Flash-001 and Pro-002 models and GPT-4o-0827 (OpenAI, 2024).
We use Wikidata5m (Wang et al., 2021) as our knowledge graph after preprocessing (check Appendix B.1.1 for details). To generate challenging and diverse specifications, we sample 50 pivot nodes from two populations: the top 2000 nodes by out-degree in the global graph, and nodes whose subgraph within radius $\rho$ contains at least 2000 vertices. This strategy ensures specifications rooted around any of the pivot nodes have a large number of paths, making enumerative certification (where all possibilities are tested for satisfaction of the specification) impractical. We set the maximum path length parameter as $\rho=5$, as we empirically observe that longer paths could result in queries that are very unrelated to the head node of the path. As our certificates are over all paths with lengths at most $\rho$ in a given subgraph, we equally prioritize the different possible path lengths in $[1,\rho]$, even though paths with longer lengths can be fewer in number than those with shorter lengths. Hence, we define our sampler from our distribution over paths (Algorithm 1, line 2) which first selects a path length from $[1,\rho]$ uniformly randomly and then selects a path having the chosen path length also uniformly randomly. A query is constructed from the selected path by uniformly selecting any aliases for the nodes in the path (Algorithm 1, line 3). We use uniform distributions so as not to bias the results by prioritizing some elements of the underlying sample spaces. Note, however, the distributions can be modified according to the specific certification usecases. We study the variation in the performance of LLMs with varying path lengths within $[1,\rho]$ in Appendix A.
A prompt is constructed for a selected path and query by prepending a context containing the relevant information to answer the query. The design of the context results in $3$ kinds of specifications, which we describe in Section 4.2. We provide some few-shot examples in the context as well, to explain the task to the models. We provide an ablation study on the effects of varying the number of few-shot examples on the performance of the target LLM in Appendix A. We formulate our prompts as multiple choice questions to facilitate evaluation of the LLM responses. The distribution of prompts is primarily determined by the query generation and answer option selection processes, with occasional context adjustments to accommodate model-specific context window constraints. Queries are derived from paths with uniformly distributed lengths, ensuring a balanced representation of reasoning complexity. Answer options are sampled non-uniformly, prioritizing distractors (in the distractor setting), followed by entities from the query path, and finally random entities related to path nodes. This approach balances query complexity and creates challenging answer options. We detail the complete prompt construction process in Appendix B.3. For evaluation of LLM responses, we use a string matching procedure. The detailed design of our evaluator is presented in Appendix B.7. QuaCer-C generates certificates with confidence $1-\delta=95\%$ and number of samples, $n=250$ samples. We conduct our experiments for open-source LLMs on 2 A100 GPUs.
4.2 Different kinds of specifications
We study the certificates for $3$ different kinds of specifications that arise from variations in the construction of context in the prompts to LLMs (Algorithm 1, line 4) — with context shuffling and distractor context (Shuffle Distractor), with context shuffling and without distractor context (Shuffle), and without context shuffling and without distractor context (Vanilla). These settings enable us to study the effects of these operations on the knowledge provided in LLM prompts. When distractor context is provided, it is only for $1$ distractor node, so as to fit the relevant context for the nodes on the considered path within the context windows of the target LLMs. We hypothesize that distractors to nodes later in the path, closer to the tail node which consists of the final answer, would be more challenging for the LLM due to their proximity and similarity to the answer node. Our distractor node sampler (Appendix B.3, Algorithm 4) thus employs a weighted sampling approach to prioritize distractor nodes closer to the path’s tail. We provide an ablation study on the effects of varying the distribution of distractor nodes in Appendix A.
Table 1: Certification results for different LLMs
Model Specification kind Precision Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Pro Vanilla - $0.72\pm 0.06$ $0.83\pm 0.05$ $0.78\pm 0.06$
Shuffle - $0.71\pm 0.06$ $0.82\pm 0.05$ $0.77\pm 0.06$
Shuffle Distractor - $0.64\pm 0.09$ $0.75\pm 0.08$ $0.70\pm 0.09$
GPT-4o Vanilla - $0.70\pm 0.06$ $0.81\pm 0.06$ $0.76\pm 0.06$
Shuffle - $0.69\pm 0.06$ $0.80\pm 0.06$ $0.75\pm 0.06$
Shuffle Distractor - $0.62\pm 0.09$ $0.74\pm 0.08$ $0.68\pm 0.09$
Gemini-1.5-Flash Vanilla - $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Shuffle - $0.45\pm 0.06$ $0.57\pm 0.06$ $0.51\pm 0.06$
Shuffle Distractor - $0.42\pm 0.10$ $0.55\pm 0.10$ $0.48\pm 0.10$
Llama (8B) Vanilla fp16 $0.31\pm 0.09$ $0.43\pm 0.10$ $0.36\pm 0.10$
8bit $0.30\pm 0.05$ $0.42\pm 0.06$ $0.36\pm 0.06$
4bit $0.27\pm 0.05$ $0.39\pm 0.05$ $0.33\pm 0.05$
Shuffle fp16 $0.31\pm 0.05$ $0.44\pm 0.06$ $0.37\pm 0.06$
8bit $0.31\pm 0.06$ $0.44\pm 0.06$ $0.37\pm 0.06$
4bit $0.28\pm 0.07$ $0.40\pm 0.07$ $0.34\pm 0.07$
Shuffle Distractor fp16 $0.30\pm 0.10$ $0.42\pm 0.11$ $0.36\pm 0.11$
8bit $0.30\pm 0.06$ $0.42\pm 0.06$ $0.35\pm 0.06$
4bit $0.25\pm 0.09$ $0.36\pm 0.09$ $0.30\pm 0.09$
Mistral (7B) Vanilla fp16 $0.33\pm 0.05$ $0.45\pm 0.05$ $0.39\pm 0.05$
8bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.05$
4bit $0.31\pm 0.06$ $0.43\pm 0.06$ $0.37\pm 0.06$
Shuffle fp16 $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
8bit $0.34\pm 0.06$ $0.46\pm 0.06$ $0.40\pm 0.06$
4bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
Shuffle Distractor fp16 $0.33\pm 0.05$ $0.45\pm 0.05$ $0.39\pm 0.05$
8bit $0.33\pm 0.11$ $0.46\pm 0.12$ $0.39\pm 0.12$
4bit $0.28\pm 0.11$ $0.39\pm 0.12$ $0.33\pm 0.11$
Phi-3 (14B) Vanilla fp16 $0.35\pm 0.05$ $0.47\pm 0.05$ $0.41\pm 0.05$
8bit $0.35\pm 0.05$ $0.47\pm 0.05$ $0.41\pm 0.05$
4bit $0.33\pm 0.04$ $0.45\pm 0.05$ $0.39\pm 0.05$
Shuffle fp16 $0.35\pm 0.04$ $0.48\pm 0.04$ $0.41\pm 0.04$
8bit $0.34\pm 0.06$ $0.47\pm 0.06$ $0.40\pm 0.06$
4bit $0.33\pm 0.04$ $0.46\pm 0.05$ $0.39\pm 0.05$
Shuffle Distractor fp16 $0.33\pm 0.11$ $0.45\pm 0.11$ $0.38\pm 0.11$
8bit $0.31\pm 0.08$ $0.43\pm 0.09$ $0.37\pm 0.08$
4bit $0.30\pm 0.08$ $0.42\pm 0.09$ $0.36\pm 0.09$
Phi-3 (3B) Vanilla fp16 $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
8bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
4bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
Shuffle fp16 $0.34\pm 0.04$ $0.47\pm 0.05$ $0.40\pm 0.05$
8bit $0.32\pm 0.04$ $0.44\pm 0.05$ $0.38\pm 0.05$
4bit $0.32\pm 0.05$ $0.44\pm 0.05$ $0.38\pm 0.05$
Shuffle Distractor fp16 $0.32\pm 0.10$ $0.45\pm 0.10$ $0.38\pm 0.10$
8bit $0.31\pm 0.09$ $0.43\pm 0.10$ $0.37\pm 0.10$
4bit $0.29\pm 0.10$ $0.41\pm 0.10$ $0.35\pm 0.10$

### v3

@@ 2.1.1 Natural noise in prompts
(Distractor). Consider DAG ${\mathcal{G}}^{\prime}$ in ${\mathcal{G}}$, with nodes ${\mathcal{N}}^{\prime}$ and ${\mathcal{N}}$ respectively. A distractor node $\tilde{v}\notin{\mathcal{N}}^{\prime}$ satisfies $\exists v\in{\mathcal{N}}\setminus{\mathcal{N}}^{\prime},((v,\tilde{v})\in{\mathcal{E}})\wedge(\exists v^{\prime}\in{\mathcal{N}}^{\prime},(v,\tilde{v})_{\mathcal{A}}=(v,v^{\prime})_{\mathcal{A}}\wedge v^{\prime}\neq{\mathcal{N}}^{\prime}[-1])$.
Shuffling. Prior works (Chen et al., 2024) have shown that LM performance can vary with information ordering. Hence, we shuffle the information in provided in prompt, to specify that LM’s response should be invariant to information ordering.
@@
Our property quantifies the probability of observing correct knowledge comprehension for a random multi-hop reasoning prompt, with possible natural noise, developed from ${\mathcal{G}}$. We formally define the property for ${\mathcal{L}}$ as a probabilistic program over ${\mathcal{G}}$ in Alg 1. We follow the syntax of the imperative probabilistic programming language in (Sankaranarayanan et al., 2013, Figure 3). The language has primitives for sampling from common distributions like Uniform ($\mathcal{U}$), Bernoulli ($Ber$), etc., and estimateProb(.) function that outputs the probability of a random variable attaining a certain value. We use a generic identifier $\mathcal{D}$ (line 1) for samplers of discrete distributions ($\dots$’ denotes samplers for other discrete distributions). We use any(.) function to denote that at least $1$ of its inputs is true.
A prompt for ${\mathcal{L}}$ consists of ${\mathcal{Q}}$ and a context $\Gamma$ containing information relevant to answer ${\mathcal{Q}}$ (Alg 1, line 6). ${\mathcal{Q}}$ is developed from a randomly sampled DAG ${\mathcal{G}}^{\prime}$, given by the function sampleDAG. We elaborate on sampleDAG() in §2.1.2. ${\mathcal{G}}^{\prime}$ is transformed into a natural language query with a randomly-selected template from a set of prespecified templates, $\tau$. The template samples random aliases of nodes in ${\mathcal{N}}^{\prime}_{source}$ and relations in ${\mathcal{G}}^{\prime}$ to form a natural language query for ${\mathcal{N}}^{\prime}_{sink}$. $\Gamma$ is formed by concatenating ($\odot$) contexts for all nodes in ${\mathcal{N}}^{\prime}$, followed by optional shuffling. We denote distractor text in $\Gamma$ as context of randomly sampled nodes from a distribution $\mathcal{D}$ over all distractor nodes of ${\mathcal{N}}^{\prime}$ in ${\mathcal{N}}$. Our final certificate (line 7) quantifies the probability that ${\mathcal{L}}$ generates any alias of ${\mathcal{N}}^{\prime}_{sink}$, which is correct answer. The certificate depends on the different sampling steps and is specific to ${\mathcal{G}}$.
Input: ${\mathcal{L}},{\mathcal{G}},\tau,\texttt{args}$
@@
2:  ${\mathcal{G}}^{\prime}\coloneqq\texttt{sampleDAG}({\mathcal{G}},\texttt{args})$
3:  ${\mathcal{N}}^{\prime}\coloneqq\texttt{topologicalOrder}({\mathcal{G}}^{\prime})$
4:  ${\mathcal{Q}}\coloneqq\mathcal{D}(\tau)({\mathcal{G}}^{\prime},{\mathcal{N}}^{\prime})$
5:  $\Gamma\coloneqq\texttt{shuffle}([{\mathcal{N}}^{\prime}[0]_{\gamma},\dots,{\mathcal{N}}^{\prime}[-1]_{\gamma},(\mathcal{D}({\mathcal{N}}))_{\gamma},\dots])$
6:  $\mathcal{P}\coloneqq\Gamma\odot{\mathcal{Q}}$
7:  $p\coloneqq\texttt{estimateProb}(\texttt{any}({\mathcal{L}}(\mathcal{P})=={\mathcal{N}}^{\prime}[-1]_{\mathcal{A}}))$
Algorithm 1 General certification
Input: ${\mathcal{G}},\rho$; Output: ${\mathcal{G}}^{\prime}$
1:  $l\coloneqq\mathcal{D}(\{2,\dots,\rho\})$
2:  $\mathcal{B}\coloneqq\texttt{boundedDAGs}({\mathcal{G}},{\mathcal{N}}_{source},l)$
3:  ${\mathcal{G}}^{\prime}\coloneqq\mathcal{D}(\mathcal{B})$
Algorithm 2 Entity centric sampleDAG
Input: ${\mathcal{G}},{\mathcal{G}}_{ref}$; Output: ${\mathcal{G}}^{\prime}$
1:  $\mathcal{I}\coloneqq\texttt{generateIsomorphisms}({\mathcal{G}},{\mathcal{G}}_{ref})$
2:  ${\mathcal{G}}^{\prime}\coloneqq\mathcal{D}(\mathcal{I})$
Algorithm 3 Relations centric sampleDAG
2.1.2 Local specifications
We define individual knowledge comprehension specifications on queries over DAGs that share a common characteristic, thus forming local specifications (Seshia et al., 2018). Local specifications are commonly defined for and used in neural network verification (Singh et al., 2019; Baluta et al., 2021), due to their tractability to verify and simplicity to interpret. Although there can be various common characteristics to specify DAGs, we define two kinds of local specifications with different common characteristics. sampleDAG() randomly samples DAGs from given ${\mathcal{G}}$ having the common characteristics.
[…]
Relations centric. Question-answering datasets are commonly formed with question templates (Cui et al., 2016) each having fixed relations (relation-centric) between variable entities of specific types. For example, “Which drug is a treatment for \${disease}?”, where \${disease} can be substituted with any curable disease to form individual questions. We define specifications by fixing the structure and relations of DAG and randomly sampling nodes to obtain the final DAG over which (relation-centric) queries are formed. Alg 3 defines the sampling procedure for ${\mathcal{G}}^{\prime}$ to obtain relation-centric DAGs from reference ${\mathcal{G}}_{ref}$.
@@ 2.1.2 Local specifications
(Isomorphic DAGs) Let finite DAGs ${\mathcal{G}}_{1}=({\mathcal{N}}_{1},{\mathcal{E}}_{1})$ and ${\mathcal{G}}_{2}=({\mathcal{N}}_{2},{\mathcal{E}}_{2})$ be isomorphic when there exists a bijection $\phi$ between the nodes of ${\mathcal{G}}_{1}$ and ${\mathcal{G}}_{2}$ such that the following conditions hold. (1) $\forall v_{1},v_{2}\in{\mathcal{N}}_{1},(v_{1},v_{2})\in{\mathcal{E}}_{1}\Leftrightarrow(\phi(v_{1}),\phi(v_{2}))\in{\mathcal{E}}_{2}$; (2) $\forall(v_{1},v_{2})\in{\mathcal{E}}_{1},(v_{1},v_{2})_{\mathcal{A}}=(\phi(v_{1}),\phi(v_{2}))_{\mathcal{A}}$. generateIsomorphisms() gives all isomorphisms of ${\mathcal{G}}_{ref}$ from ${\mathcal{G}}$.
@@ 2.2 Certification method
Our algorithm certifies the target LLM ${\mathcal{L}}$ by computing an interval $[p_{l},p_{u}]$ bounding probability $p$ (Alg 1, line 7) for a given specification distribution over ${\mathcal{G}}$ with high confidence. We model $p$ as the probability of setting the underlying boolean random variable $\mathcal{R}\triangleq({\texttt{any}({\mathcal{L}}(\mathcal{P})=={\mathcal{N}}^{\prime}[-1]_{\mathcal{A}})})$ to true (success). Thus, $\mathcal{R}\sim Ber(p)$. Exactly determining $p$ would require enumerating over all possible $\mathcal{P}$ which can be developed from any DAG of ${\mathcal{G}}$ with any random aliases, resulting in an infeasible number of possible prompts, as shown in Appendix D.8. Moreover, we want our method to generalize to closed-source LLMs as well, where the internal structures of the models are unknown. Hence, we cannot use any symbolic methods (Mirman et al., 2020) to determine $p$. Thus, to scalably certify the black-box target LM ${\mathcal{L}}$, we estimate $p$ with provably high-confidence (low error) bounds. Confidence is defined as the probability of the true value $p$ being within the bounds, i.e., $Pr[p\in[p_{l},p_{u}]]$. To establish formal guarantees, we want our estimation procedure to be such that the actual confidence is at least the user-specified confidence level, $1-\delta$ (i.e., $Pr[p_{l}\leq p\leq p_{u}]\geq 1-\delta$), where $\delta>0$ is a small constant. Hence, we use Clopper-Pearson (CP) confidence intervals (Clopper & Pearson, 1934; Kurz et al., 2014), which is a conservative method known to produce guaranteed high confidence intervals. We show empirical evidence of the bounds’ conservativeness in Appendix C. To compute high-confidence bounds on $p$, we make $n$ independent and identically distributed observations of $\mathcal{R}$, in which we obtain $k$ successes, $k\in[0,n]$. We generate CP confidence intervals with the $n$ observations to bound $p$ with $1-\delta$ confidence.
@@ 3 Experiments
Our experiments were run on 2 A100 GPUs with 40GB VRAM each. We certify the following open-source models — Llama-3.2-instruct 1B, 3B, 11B (Dubey et al., 2024), Phi-3 3B and 14B (abdin2024Phi-3technicalreporthighly) and their 4-bit quantizations to study the effects of quantization on knowledge comprehension. Among closed-source models with API access, we certify Gemini-1.5 (Gemini Team, 2024) Flash-001 and Pro-002 and GPT-4o-0827 (OpenAI, 2024). We use PrimeKG (Chandak et al., 2023) and Wikidata5m (Wang et al., 2021b) as our knowledge graphs (KGs) (details in Appendix D.3.1 and Appendix D.2.1 respectively). PrimeKG is a KG for precision medicine, while Wikidata5m is for general knowledge from Wikipedia. These practical KGs are chosen only for illustration, and our framework generalizes to other KGs. We show relation-centric and entity-centric specifications for PrimeKG and Wikidata5,m respectively.
PrimeKG. We manually define 9 relation-centric specifications based on common clinical reasoning patterns. All specifications are shown in §D.4.2. Each specification is defined by: 1. A reference DAG (${\mathcal{G}}_{ref}$); 2. A set of natural language templates (5-7 per specification); 3. Constraints on entity types for isomorphic DAG sampling.
An example specification is the “drug-disease interaction” specification. It consists of ${\mathcal{G}}_{ref}$ which is a 3-node DAG (shown on the right) with drug→disease edges. An example template to form natural language queries from ${\mathcal{G}}_{ref}$ is “Which drug treats disease B but is contraindicated for disease A?”.
Wikidata5m. We created 50 entity-centric specifications using Wikidata5m, sampling linear paths (DAGs) up to length $\rho=5$ to maintain query relevance. The 50 starting nodes were selected from high out-degree (top 2k) or high-density neighborhood ($>$2k nodes within radius $\rho$) populations the specifications have a large number of paths, making enumeration infeasible. Note that these specifications are only for demonstration and LLMCert-C is not limited to them. To avoid bias towards the naturally more frequent shorter paths, we first sample a desired path length uniformly from $[2,\rho]$. The actual path is then generated from the starting node according to this length (Alg. 2). We analyze the impact of varying query path lengths on LLM performance in Appendix E. Note that the framework is adaptable to other distributions as needed by the specific certification usecase.
Specifications with varying complexities. We certify $2$ kinds of specifications with different kinds of contexts in LLM prompts (Alg 1, line 4) — with context shuffling and distractor context (Distractor) and without such noise (Vanilla). This enables us to study the effects of natural noise on LLM performance. Leveraging the linear structure of Wikidata5m DAGs, we use weighted sampling (Alg. 6) to select distractors closer to the answer node, hypothesizing this increases difficulty. Since PrimeKG lacks such unique linearizations, its distractor nodes are sampled uniformly. We provide an ablation study on the effects of varying the weighted distractor sampling in Appendix E. Prompts give multiple answer options with $1$ correct, to simply evaluate LLM responses using string matching (Appendix D.9). For answer options in Distractor setting, we prioritize distractor nodes, followed by nodes from the query path, and finally, random nodes related to path nodes. Vanilla setting is similar, without distractors. All further details of prompt construction are in Appendix D.5.
Baseline. We compare LLMCert-C’s results with a benchmarking baseline. It measures the target LLM’s accuracy on a static dataset. We generate 50 static queries from each specification by uniformly sampling DAGs. Queries use a fixed alias of entities and relations, with no shuffling or distractors in contexts to remove natural noise.
Table 1: Average certification bounds for specifications over Wikidata5m and PrimeKG KGs for different LLMs. Bounds are given as [lower bound, upper bound].
Model Quantization Wikidata PrimeKG
Baseline Vanilla Distractor Baseline Vanilla Distractor
Gemini-1.5-Pro - $0.87$ [0.80, 0.89] [0.64, 0.75] $1.00$ [0.97, 1.00] [0.92, 0.98]
GPT-4o - $0.88$ [0.80, 0.89] [0.62, 0.74] $0.99$ [0.97, 1.00] [0.85, 0.95]
Gemini-1.5-Flash - $0.81$ [0.72, 0.83] [0.43, 0.56] $0.99$ [0.96, 0.99] [0.89, 0.96]
Phi-3 (14B) fp16 $0.63$ [0.54, 0.66] [0.33, 0.45] $0.98$ [0.93, 0.98] [0.75, 0.85]
4bit $0.63$ [0.52, 0.65] [0.30, 0.42] $0.95$ [0.92, 0.97] [0.71, 0.81]
Llama (11B) fp16 $0.61$ [0.50, 0.63] [0.33, 0.45] $0.97$ [0.91, 0.97] [0.84, 0.92]
4bit $0.54$ [0.45, 0.57] [0.29, 0.41] $0.94$ [0.88, 0.95] [0.79, 0.88]
Phi-3 (3B) fp16 $0.58$ [0.50, 0.61] [0.32, 0.45] $0.92$ [0.85, 0.92] [0.70, 0.81]
4bit $0.57$ [0.48, 0.60] [0.29, 0.41] $0.89$ [0.83, 0.91] [0.74, 0.83]
Llama (3B) fp16 $0.51$ [0.43, 0.55] [0.27, 0.39] $0.87$ [0.80, 0.89] [0.70, 0.80]
4bit $0.46$ [0.39, 0.51] [0.24, 0.35] $0.84$ [0.73, 0.83] [0.63, 0.73]
Llama (1B) fp16 $0.38$ [0.30, 0.41] [0.22, 0.33] $0.48$ [0.43, 0.56] [0.39, 0.52]
4bit $0.34$ [0.27, 0.38] [0.20, 0.31] $0.38$ [0.33, 0.45] [0.28, 0.40]